OpenStudy (anonymous):

The one-to-one function h is defined by h(x)= 9x-8/3+8x . Find h^-1, the inverse of h. Then, give the domain and range of h^-1 using interval notation.

OpenStudy (anonymous):

please and thank you =)

OpenStudy (anonymous):

let y = 9x-8/3+8x then 3y+8xy=9x-8 x(8y-9)=-3y-8 x= 3y+8/9-8y h^-(x) = 3y+8/9-8y

OpenStudy (anonymous):

i dont understand the domain and range

OpenStudy (anonymous):

domain is all real numbers except 9/8 range all real numbers

OpenStudy (anonymous):

would it be written with open or closed paranthesies

OpenStudy (anonymous):

the answer asks for open or closed lol

OpenStudy (anonymous):

open we cannot include infinity

OpenStudy (anonymous):

im assuming the range is both open

OpenStudy (anonymous):

for the domain are any closed so would it be??? (-infinity,8) (8,9) (9,infinity)

OpenStudy (anonymous):

i think i got the range..just confused on the domain

OpenStudy (anonymous):

domain is set for all function is defined and have finite value so R-{9/8}

OpenStudy (anonymous):

the range is (-infinity,infinity)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol sorry just trying to understand this..i appreciate your help

OpenStudy (anonymous):

the domain is the same as the range?

OpenStudy (anonymous):

no in domain we cannot include the value 9/8 because for this value denominator becomes 0 and function not defined

OpenStudy (anonymous):

so how i wrote it was right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!!!!!!!